Solve the formula v=1/3ah for h
Answer by nyc_function(2741)

To solve for h means to isolate h.

V = 1/3 * ah, which can be also written V = ah/3.

Multiply both sides by 3.

3V = (ah/3) * 3

3V = ah

We now divide both sides by a.

3V/a = h

Done!